# L-equivalence and Fourier–Mukai partners of cubic fourfolds

### v2

@@ 1. Introduction
We prove Proposition 1.5 by constructing a Counting Formula for Fourier–Mukai partners of certain cubic fourfolds. Specifically, we make the following simplifying assumption:
Assumption 1.7 (Assumption 3.1).
There is an equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$ if and only if there is a Hodge isometry $T(X)\simeq T(Y)$.
This assumption is satisfied for a large class of cubic fourfolds (see Proposition 3.3). Inspired by the works [19, 34, 9], we prove the following:
@@
Let $X$ be a cubic fourfold satisfying Assumption 1.7. Assume moreover that $\operatorname{disc}(T(X))$ is not divisible by $3$. Then we have
@@ 6. LF:
For two cubic fourfolds $X$, $Y$, we have the following implications, where the implications marked by $\dagger$ are proved only when $X$ and $Y$ satisfy Assumption 1.7 and $X$ satisfies $\operatorname{End}_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}$:
@@
The fact that $\mathbb{L}$-equivalent cubic fourfolds are Fourier–Mukai partners was recently independently proved in [4] for very general cubic fourfolds in $\mathcal{C}_{d}$ for $d\equiv 0,2\pmod{6}$ not divisible by $9$. We note that such cubic fourfolds satisfy Assumption 1.7. They also prove LE $\implies$ LF and the fact that BE does not imply LE.
For the convenience of the reader, let us complete Diagram 1.2 with the implications that are already known. The implication FM $\implies$ DF was proved recently in [29], and the fact that BF $\implies$ DF follows from the proof of the $D$-equivalence conjecture for hyperkähler manifolds of $\operatorname{K3}^{[n]}$-type of [37]. Finally, under Assumption 1.7, we also have DF $\implies$ FM by [3, Corollary 9.3]. The question of whether implication FM $\implies$ BE holds is a conjecture by Huybrechts [24, Conjecture 7.3.21], and the implication BF $\implies$ BE is conjectured in [8, Conjecture 1.2]. For a survey on these equivalence relations, we refer to [8].
@@ 2.1. Lattice theory
[44, Proposition 1.4.2] Let $f\colon N\to L$ and $g\colon N\to L^{\prime}$ be two even overlattices of an even lattice $N$.
@@
[44, Proposition 1.5.1] Let $N$ be an even lattice and let $f\colon N\hookrightarrow L$ and $g\colon N\hookrightarrow L^{\prime}$ be primitive embeddings into even lattices $L$, $L^{\prime}$. Write $T\coloneqq N^{\perp}\subset L$ and $T^{\prime}\coloneqq N^{\perp}\subset L^{\prime}$.
@@ 2.3. Hodge realisation
where the (1)(-1) indicates that we rescaled the bilinear forms by 1-1. Since the divisibility of $\lambda_{1}$ in $\tilde{H}(\mathcal{A}_{X},\mathbb{Z})$ is $1$, it follows that the coarseness of $F(X)$ is $\operatorname{crs}(F(X))=1$ (see [40, Definition 2.4] for the definition of coarseness). Therefore, by [40, Theorem 3.5], we find
@@
Suppose that there is a K3 surface $S$ with a Hodge isometry $T(X)(-1)\simeq T(S)$, so that $F(X)$ is a moduli space of stable sheaves on $S$ by [3, Corollary 9.6]. Then Lemma 2.19 is equivalent to saying that $F(X)$ is a fine moduli space of sheaves on $S$ by [38, Theorem 3.9].
@@
If $X$ and $Y$ are K3 surfaces, the statement is follows from [39, Proposition 3.11] combined with [11, Lemma 3.7].
If $X$ and $Y$ are cubic fourfolds, we have $[T(X)(1)]=[T(Y)(1)]$ in $K_{0}^{\oplus}(\operatorname{HS}_{\mathbb{Z},2})$. Since $\operatorname{End}_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}$, this implies by [11, Lemma 3.7] that there exists an isomorphism of Hodge structures (not necessarily a Hodge isometry) $T(X)(1)\simeq T(Y)(1)$. Moreover, since $X$ and $Y$ are $\mathbb{L}$-equivalent, we have $\operatorname{disc}(T(X))=\operatorname{disc}(T(Y))$ by Lemma 2.17 and Lemma 2.18. It now follows from [39, Corollary 3.10] that we have a Hodge isometry $T(X)(1)\simeq T(Y)(1)$, which clearly induces a Hodge isometry $T(X)\simeq T(Y)$ via the Tate twist. ∎
@@ 3. Counting Fourier–Mukai partners of cubic fourfolds
In this section, we consider a cubic fourfold $X$ which satisfies the following Derived Torelli assumption:
Assumption 3.1.
For any cubic fourfold $Y$, the following two statements are equivalent:
  1. (1)
  2. (2)
@@
Notice that the implication $(1)\Rightarrow(2)$ always holds. By [23, Proposition 3.4], any equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$ induces an orientation-preserving Hodge isometry $\widetilde{H}(\mathcal{A}_{X},\mathbb{Z})\to\widetilde{H}(\mathcal{A}_{Y},\mathbb{Z})$ ([23, Theorem 1.2]). Then, by [2, Proposition 2.3], there is an identification $H^{4}_{\mathrm{prim}}(X,\mathbb{Z})\cong A_{2}^{\perp}\subset\widetilde{H}(\mathcal{A}_{X},\mathbb{Z})$. By Voisin’s proof of the integral Hodge conjecture for cubic fourfolds, the primitive algebraic lattices coincide with the primitive Hodge classes, $A(X)_{\mathrm{prim}}=H^{4}_{\mathrm{prim}}(X,\mathbb{Z})\cap H^{2,2}(X)$. Hence we have a Hodge isometry $T(X)=A(X)_{\mathrm{prim}}^{\perp}\cong A(Y)_{\mathrm{prim}}^{\perp}=T(Y)$.
@@
In each of the following cases, Assumption 3.1 holds:
@@
Let $X$ be a cubic fourfold satisfying Assumption 3.1. Assume moreover that $\operatorname{disc}(T(X))$ is not divisible by $3$. Then we have
@@
Let $X$ be a cubic fourfold satisfying Assumption 3.1. Let $N^{\prime}$ be a lattice. Then there is a one-to-one correspondence between the set $\operatorname{FM}(X,N^{\prime})$ and the set of $O_{\mathrm{Hodge}}(T(X))$-orbits of primitive embeddings in $\operatorname{Emb}_{N^{\prime}}(T(X),\Lambda_{\text{cub}}^{0})$.
@@
For surjectivity, note that any primitive embedding $i\colon T(X)\hookrightarrow\Lambda_{\text{cub}}^{0}$ induces a Hodge structure on $\Lambda_{\text{cub}}^{0}$. We denote the resulting Hodge lattice $L_{i}$, i.e. $L_{i}$ is the Hodge lattice whose underlying lattice structure is $\Lambda_{\text{cub}}^{0}$ and whose Hodge structure is induced by $i$. Since $T(X)^{\perp}\simeq N^{\prime}\in\mathcal{H}(N)$, it follows that $N^{\prime}$ contains no vectors of square $6$ and divisibility $3$, hence there is a unique cubic fourfold $Y$ for which there is a Hodge isometry $H^{4}(Y,\mathbb{Z})_{\text{prim}}\simeq L_{i}$. By construction, we have $T(Y)\simeq T(L_{i}(1))\simeq T(X)$, and $A(Y)_{\text{prim}}\simeq\operatorname{NS}(L_{i}(1))\simeq N^{\prime}$, where $T(L_{i}(1))$ and $\operatorname{NS}(L_{i}(1))$ are defined as in Section 2.3. Since $X$ satisfies Assumption 3.1, it follows that $Y\in\operatorname{FM}(X,N^{\prime})$. Clearly, we have $i=\operatorname{emb}(Y)$. Therefore, the map $\operatorname{emb}$ is surjective, as required. ∎
@@
Let $X$ be a cubic fourfold satisfying Assumption 3.1, and such that $\operatorname{disc}(T(X))$ is not divisible by $3$. Then we have
@@
Let $X$ be a cubic fourfold satisfying Assumption 3.1, and write $N\coloneqq A(X)_{\mathrm{prim}}$. Suppose that $\operatorname{disc}(N)$ is not divisible by $3$. Then for any $N^{\prime}\in\mathcal{H}(N)$, we have
@@ 4. Examples of $\mathbb{L}$-equivalent cubic fourfolds
[25, §5] [30, §2.6.1] Let $X$ be a cubic fourfold with one of the following properties:
@@
In light of Proposition 4.1, Pfaffian cubic fourfolds seem like a good place to start looking for $\mathbb{L}$-equivalent cubic fourfolds. However, very general Pfaffian cubic fourfolds have no Fourier–Mukai partners. This result follows immediately from [13, Proposition 2.6], and also from [50, Theorem 1.1] combined with [45, Proposition 1.10]. Here, we give a proof using Theorem 3.5.
@@ 5. Implications of $\mathbb{L}$-equivalence
The following proposition is the analogue for cubic fourfolds of [11, Proposition 3.3]. We remark that under Assumption 3.1 this follows immediately from [21, Theorem 1.1].
@@
The following theorem is the analogue for cubic fourfolds of [11, Theorem 3.4]. Again, under Assumption 3.1, Theorem 5.4 implies that $\mathbb{L}$-equivalent cubic fourfolds are Fourier-Mukai partners, and the results again follows from [21, Theorem 1.1].
@@
Let $X$ and $Y$ be cubic fourfolds satisfying Assumption 3.1 (for example, if $X$ is a cubic fourfold satisfying one of the properties of Proposition 3.3) Moreover, assume $\operatorname{rk}(T(X))\neq 4$ and $\operatorname{End}_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}$. Then, if $X$ and $Y$ are $\mathbb{L}$-equivalent, they are Fourier–Mukai partners.
@@
By Theorem 2.20, there is a Hodge isometry $T(X)\cong T(Y)$. Then, by Assumption 3.1, $X$ and $Y$ are Fourier–Mukai partners. ∎
@@
Since $X$ and $Y$ are assumed to be $\mathbb{L}$-equivalent, $X^{[2]}$ and $Y^{[2]}$ are $\mathbb{L}$-equivalent by [46, Lemma 2.1]. Now the result follows directly from the beautiful formula of Galkin–Shinder [14]: for any cubic fourfold $X$, we have
@@
Let $X$ and $Y$ be cubic fourfolds. Assume that $X$ satisfies Assumption 3.1, that $\operatorname{rk}(T(X))\neq 4$, and that $\operatorname{End}_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}$. Then if $F(X)\sim_{\mathbb{L}}F(Y)$, there is an equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$, hence also an equivalence $\mathcal{D}^{b}(F(X))\simeq\mathcal{D}^{b}(F(Y))$.
@@
By [40, Theorem 3.5] and Lemma 2.19, the gluing group
@@
has order $|G|=\operatorname{disc}(T(F(X)))$. Therefore, if $F(X)$ is $\mathbb{L}$-equivalent to $F(Y)$, it follows from Lemma 2.17 that $\operatorname{disc}(T(F(X)))=\operatorname{disc}(T(F(Y)))$. Since $T(F(X))$ is Hodge anti-isometric to $T(X)$, we find that $[T(X)(1)]=[T(Y)(1)]$. By Theorem 2.20, there is a Hodge isometry $T(X)\simeq T(Y)$, hence an equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$ by Assumption 3.1. The fact that there is an equivalence $\mathcal{D}^{b}(F(X))\simeq\mathcal{D}^{b}(F(Y))$ follows from the main result of [29]. ∎
@@
Now apply the motivic volume morphism $\operatorname{Vol}_{K}:K_{0}(\operatorname{Var}_{K})\to K_{0}^{\hat{\mu}}(\operatorname{Var}_{\mathbb{C}})$ of [43, Theorem 3.1.1]. From $\mathcal{X}_{K}]=[\mathcal{Y}_{K}]$ we get $\operatorname{Vol}_{K}([\mathcal{X}_{K}])=\operatorname{Vol}_{K}([\mathcal{Y}_{K}])$. From the fact that $\mathcal{X}_{R}$ ($\mathcal{Y}_{R}$) is a smooth proper $R$ model of $\mathcal{X}_{K}$ ($\mathcal{Y}_{K}$), so $\operatorname{Vol}_{K}([\mathcal{X}_{K}])=[\mathcal{X}_{0}]=[X]$ and $\operatorname{Vol}_{K}([\mathcal{Y}_{K}])=[\mathcal{Y}_{0}]=[Y]$. This concludes the proof. ∎

### v3

@@ 1. Introduction
We prove Proposition 1.5 by constructing a Counting Formula for Fourier–Mukai partners of certain cubic fourfolds. Specifically, we will work with cubic fourfolds satisfying the following Definition.
Definition 1.7.
A cubic fourfold $X$ is said to satisfy Derived Torelli if it has the property that for any cubic fourfold $Y$, the following two statements are equivalent:
  1. (1)
  2. (2)
Note that $(1)\implies(2)$ always holds, see Remark 3.1. There is a large class of cubic fourfolds which satisfy Derived Torelli, for instance if $X$ is a general cubic fourfold not contained in a Hassett divisor (see also Proposition 3.2 and Remark 3.3).
Inspired by the works [19, 35, 9], we prove the following:
@@
Let $X$ be a cubic fourfold which satisfies Derived Torelli. Assume moreover that $\operatorname{disc}(T(X))$ is not divisible by $3$. Then we have
@@ 6. LF:
For two cubic fourfolds $X$, $Y$, we have the following implications, where the implications marked by $\dagger$ are proved only when $X$ and $Y$ satisfies Derived Torelli and $X$ is such that $\operatorname{End}_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}$:
@@
The fact that $\mathbb{L}$-equivalent cubic fourfolds are Fourier–Mukai partners was recently independently proved in [4] for very general cubic fourfolds in $\mathcal{C}_{d}$ for $d\equiv 0,2\pmod{6}$ not divisible by $9$. We note that such cubic fourfolds satisfy Derived Torelli. They also prove LE $\implies$ LF and the fact that BE does not imply LE.
For the convenience of the reader, let us complete Diagram 1.2 with the implications that are already known. The implication FM $\implies$ DF was proved recently in [30], and the fact that BF $\implies$ DF follows from the proof of the $D$-equivalence conjecture for hyperkähler manifolds of $\operatorname{K3}^{[n]}$-type of [38]. Finally, if the cubic fourfolds satisfy Derived Torelli, we also have DF $\implies$ FM by [3, Corollary 9.3]. The question of whether implication FM $\implies$ BE holds is a conjecture by Huybrechts [24, Conjecture 7.3.21], and the implication BF $\implies$ BE is conjectured in [8, Conjecture 1.2]. For a survey on these equivalence relations, we refer to [8].
@@ 2.1. Lattice theory
[45, Proposition 1.4.2] Let $f\colon N\to L$ and $g\colon N\to L^{\prime}$ be two even overlattices of an even lattice $N$.
@@
[45, Proposition 1.5.1] Let $N$ be an even lattice and let $f\colon N\hookrightarrow L$ and $g\colon N\hookrightarrow L^{\prime}$ be primitive embeddings into even lattices $L$, $L^{\prime}$. Write $T\coloneqq N^{\perp}\subset L$ and $T^{\prime}\coloneqq N^{\perp}\subset L^{\prime}$.
@@ 2.3. Hodge realisation
where the (1)(-1) indicates that we rescaled the bilinear forms by 1-1. Since the divisibility of $\lambda_{1}$ in $\tilde{H}(\mathcal{A}_{X},\mathbb{Z})$ is $1$, it follows that the coarseness of $F(X)$ is $\operatorname{crs}(F(X))=1$ (see [41, Definition 2.4] for the definition of coarseness). Therefore, by [41, Theorem 3.5], we find
@@
Suppose that there is a K3 surface $S$ with a Hodge isometry $T(X)(-1)\simeq T(S)$, so that $F(X)$ is a moduli space of stable sheaves on $S$ by [3, Corollary 9.6]. Then Lemma 2.19 is equivalent to saying that $F(X)$ is a fine moduli space of sheaves on $S$ by [39, Theorem 3.9].
@@
If $X$ and $Y$ are K3 surfaces, the statement is follows from [40, Proposition 3.11] combined with [11, Lemma 3.7].
If $X$ and $Y$ are cubic fourfolds, we have $[T(X)(1)]=[T(Y)(1)]$ in $K_{0}^{\oplus}(\operatorname{HS}_{\mathbb{Z},2})$. Since $\operatorname{End}_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}$, this implies by [11, Lemma 3.7] that there exists an isomorphism of Hodge structures (not necessarily a Hodge isometry) $T(X)(1)\simeq T(Y)(1)$. Moreover, since $X$ and $Y$ are $\mathbb{L}$-equivalent, we have $\operatorname{disc}(T(X))=\operatorname{disc}(T(Y))$ by Lemma 2.17 and Lemma 2.18. It now follows from [40, Corollary 3.10] that we have a Hodge isometry $T(X)(1)\simeq T(Y)(1)$, which clearly induces a Hodge isometry $T(X)\simeq T(Y)$ via the Tate twist. ∎
@@ 3. Counting Fourier–Mukai partners of cubic fourfolds
We begin by collecting a few remarks and criteria concerning cubic fourfolds that satisfy Derived Torelli in the sense of Definition 1.7, and we describe how to find an example of a cubic which doesn’t.
@@
Notice that the implication $(1)\Rightarrow(2)$ in Definition 1.7 always holds. By [23, Proposition 3.4], any equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$ induces an orientation-preserving Hodge isometry $\widetilde{H}(\mathcal{A}_{X},\mathbb{Z})\to\widetilde{H}(\mathcal{A}_{Y},\mathbb{Z})$ ([23, Theorem 1.2]). Then, by [2, Proposition 2.3], there is an identification $H^{4}_{\mathrm{prim}}(X,\mathbb{Z})\cong A_{2}^{\perp}\subset\widetilde{H}(\mathcal{A}_{X},\mathbb{Z})$. By Voisin’s proof of the integral Hodge conjecture for cubic fourfolds, the primitive algebraic lattices coincide with the primitive Hodge classes, $A(X)_{\mathrm{prim}}=H^{4}_{\mathrm{prim}}(X,\mathbb{Z})\cap H^{2,2}(X)$. Hence we have a Hodge isometry $T(X)=A(X)_{\mathrm{prim}}^{\perp}\cong A(Y)_{\mathrm{prim}}^{\perp}=T(Y)$.
@@
In each of the following cases, the cubic fourfold $X$ satisfies Derived Torelli:
@@
Remark 3.3.
Notice that a weaker Derived Torelli statement can be found in Conjecture 6.1 of [25], where the existence of an equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$ is conjectured to be equivalent to an orientation preserving Hodge isometry $\widetilde{H}(X,\mathbb{Z})\simeq\widetilde{H}(Y,\mathbb{Z})$. As Hodge isometries of $T(X)$ need not to lift to Hodge isometries of the Mukai lattice, it is reasonable to expect that some cubic fourfolds do not satisfy Derived Torelli in the sense of Definition 1.7.
@@
Let $X$ be a cubic fourfold that satisfies Derived Torelli (See Definition 1.7). Assume moreover that $\operatorname{disc}(T(X))$ is not divisible by $3$. Then we have
@@
Let $X$ be a cubic fourfold which satisfies Derived Torelli. Let $N^{\prime}$ be a lattice. Then there is a one-to-one correspondence between the set $\operatorname{FM}(X,N^{\prime})$ and the set of $O_{\mathrm{Hodge}}(T(X))$-orbits of primitive embeddings in $\operatorname{Emb}_{N^{\prime}}(T(X),\Lambda_{\text{cub}}^{0})$.
@@
For surjectivity, note that any primitive embedding $i\colon T(X)\hookrightarrow\Lambda_{\text{cub}}^{0}$ induces a Hodge structure on $\Lambda_{\text{cub}}^{0}$. We denote the resulting Hodge lattice $L_{i}$, i.e. $L_{i}$ is the Hodge lattice whose underlying lattice structure is $\Lambda_{\text{cub}}^{0}$ and whose Hodge structure is induced by $i$. Since $T(X)^{\perp}\simeq N^{\prime}\in\mathcal{H}(N)$, it follows that $N^{\prime}$ contains no vectors of square $6$ and divisibility $3$, hence there is a unique cubic fourfold $Y$ for which there is a Hodge isometry $H^{4}(Y,\mathbb{Z})_{\text{prim}}\simeq L_{i}$. By construction, we have $T(Y)\simeq T(L_{i}(1))\simeq T(X)$, and $A(Y)_{\text{prim}}\simeq\operatorname{NS}(L_{i}(1))\simeq N^{\prime}$, where $T(L_{i}(1))$ and $\operatorname{NS}(L_{i}(1))$ are defined as in Section 2.3. Since $X$ satisfies Derived Torelli, it follows that $Y\in\operatorname{FM}(X,N^{\prime})$. Clearly, we have $i=\operatorname{emb}(Y)$. Therefore, the map $\operatorname{emb}$ is surjective, as required. ∎
@@
Let $X$ be a cubic fourfold which satisfies Derived Torelli, and such that $\operatorname{disc}(T(X))$ is not divisible by $3$. Then we have
@@
Let $X$ be a cubic fourfold which satisfies Derived Torelli, and write $N\coloneqq A(X)_{\mathrm{prim}}$. Suppose that $\operatorname{disc}(N)$ is not divisible by $3$. Then for any $N^{\prime}\in\mathcal{H}(N)$, we have
@@ 4. Examples of $\mathbb{L}$-equivalent cubic fourfolds
[26, §5] [31, §2.6.1] Let $X$ be a cubic fourfold with one of the following properties:
@@
In light of Proposition 4.1, Pfaffian cubic fourfolds seem like a good place to start looking for $\mathbb{L}$-equivalent cubic fourfolds. However, very general Pfaffian cubic fourfolds have no Fourier–Mukai partners. This result follows immediately from [13, Proposition 2.6], and also from [51, Theorem 1.1] combined with [46, Proposition 1.10]. Here, we give a proof using Theorem 3.5.
@@ 5. Implications of $\mathbb{L}$-equivalence
The following proposition is the analogue for cubic fourfolds of [11, Proposition 3.3]. We remark that for cubic fourfolds which satisfy Derived Torelli, this follows immediately from [21, Theorem 1.1].
@@
The following theorem is the analogue for cubic fourfolds of [11, Theorem 3.4]. Again, for cubic fourfolds which satisfy Derived Torelli, Theorem 5.4 implies that $\mathbb{L}$-equivalent cubic fourfolds are Fourier-Mukai partners, and the results again follows from [21, Theorem 1.1].
@@
Let $X$ and $Y$ be cubic fourfolds which satisfy Derived Torelli (for example, if $X$ is a cubic fourfold satisfying one of the properties of Proposition 3.2) Moreover, assume $\operatorname{rk}(T(X))\neq 4$ and $\operatorname{End}_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}$. Then, if $X$ and $Y$ are $\mathbb{L}$-equivalent, they are Fourier–Mukai partners.
@@
By Theorem 2.20, there is a Hodge isometry $T(X)\cong T(Y)$. Then, if $X$ and $Y$ satisfy Derived Torelli, then they are Fourier–Mukai partners. ∎
@@
Since $X$ and $Y$ are assumed to be $\mathbb{L}$-equivalent, $X^{[2]}$ and $Y^{[2]}$ are $\mathbb{L}$-equivalent by [47, Lemma 2.1]. Now the result follows directly from the beautiful formula of Galkin–Shinder [14]: for any cubic fourfold $X$, we have
@@
Let $X$ and $Y$ be cubic fourfolds. Assume that $X$ satisfies Derived Torelli, that $\operatorname{rk}(T(X))\neq 4$, and that $\operatorname{End}_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}$. Then if $F(X)\sim_{\mathbb{L}}F(Y)$, there is an equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$, hence also an equivalence $\mathcal{D}^{b}(F(X))\simeq\mathcal{D}^{b}(F(Y))$.
@@
By [41, Theorem 3.5] and Lemma 2.19, the gluing group
@@
has order $|G|=\operatorname{disc}(T(F(X)))$. Therefore, if $F(X)$ is $\mathbb{L}$-equivalent to $F(Y)$, it follows from Lemma 2.17 that $\operatorname{disc}(T(F(X)))=\operatorname{disc}(T(F(Y)))$. Since $T(F(X))$ is Hodge anti-isometric to $T(X)$, we find that $[T(X)(1)]=[T(Y)(1)]$. By Theorem 2.20, there is a Hodge isometry $T(X)\simeq T(Y)$, hence an equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$ by Definition 1.7. The fact that there is an equivalence $\mathcal{D}^{b}(F(X))\simeq\mathcal{D}^{b}(F(Y))$ follows from the main result of [30]. ∎
@@
Now apply the motivic volume morphism $\operatorname{Vol}_{K}:K_{0}(\operatorname{Var}_{K})\to K_{0}^{\hat{\mu}}(\operatorname{Var}_{\mathbb{C}})$ of [44, Theorem 3.1.1]. From $\mathcal{X}_{K}]=[\mathcal{Y}_{K}]$ we get $\operatorname{Vol}_{K}([\mathcal{X}_{K}])=\operatorname{Vol}_{K}([\mathcal{Y}_{K}])$. From the fact that $\mathcal{X}_{R}$ ($\mathcal{Y}_{R}$) is a smooth proper $R$ model of $\mathcal{X}_{K}$ ($\mathcal{Y}_{K}$), so $\operatorname{Vol}_{K}([\mathcal{X}_{K}])=[\mathcal{X}_{0}]=[X]$ and $\operatorname{Vol}_{K}([\mathcal{Y}_{K}])=[\mathcal{Y}_{0}]=[Y]$. This concludes the proof. ∎